# When Geoscience Meets Foundation Models: Towards General Geoscience Artificial Intelligence System

### v4

@@ III-A1 Vanilla Transformer
Figure 5: The Transformer - model architecture [78].
III-A2 Visual Transformer
@@
Figure 6: The framework of ViT [90].
III-A3 Vision-Language Transformer
@@ III-B1 Supervised pre-training
III-B2 Self-supervised pre-training
Self-supervised learning (SSL) has become a cornerstone in the development of foundation models (FMs), enabling them to acquire rich features and representations from vast amounts of unlabeled data [114], [115], [116]. This approach, often referred to as the ”dark matter” of intelligence, circumvents the limitations of supervised learning by leveraging a ”semiautomatic” process to derive ”labels” directly from the data, significantly reducing time and resource costs [115]. SSL has achieved remarkable success across various domains of deep learning, particularly in natural language processing with models like BERT [97] and GPT-3 [117], and in computer vision with MAE [118] and DINOv2 [119]. The SSL workflow typically involves defining a pretext task based on unlabeled inputs to generate meaningful representations [114], [116], [120], [121]. These pretext tasks, which can be prediction-based, context-based, or generation-based, form the basis of the pretraining phase where the objective function is learned from unlabeled data [120]. The resulting representations are then transferred as initial weights for fine-tuning on specific downstream tasks. SSL methods can be broadly categorized into three approaches based on their pretext tasks: generative, contrastive, and predictive [114], [122]. This pretraining-finetuning paradigm has proven highly effective, allowing FMs to leverage the power of unlabeled data before being adapted to specific tasks with labeled data, often matching or surpassing purely supervised approaches in performance.
@@ III-B2 Self-supervised pre-training
GANs [137] represent a powerful approach in generative methods, consisting of two competing networks: a generator $G:Z\rightarrow X$ that synthesizes fake samples from random noise in the latent space, and a discriminator $D:X\rightarrow[0,1]$ that attempts to distinguish between real and fake samples. This adversarial setup has proven effective for tasks such as image generation and super-resolution. Notable advancements include BigGAN [123], which introduces innovations like a large-scale GAN architecture, conditional batch normalization, category conditioning, and orthogonal regularization, enabling the generation of high-quality, diverse, and category-specific images. In the domain of image enhancement, SRGAN [124] has demonstrated remarkable capabilities in super-resolution tasks, learning from a large dataset of downsampled images to recover high-resolution textures and effectively upscale low-resolution inputs. These applications showcase the potential of GANs not just for generating new images, but also for enhancing existing ones, pushing the boundaries of synthetic image generation and opening up numerous possibilities in image processing and computer vision tasks.
@@ III-B3 Hybrid pre-training
III-C Adaptation
Pre-trained GFMs can be adapted to specific tasks or datasets through fine tuning, prompt tuning, adapter tuning, parameter tuning, and remapping tuning. For more tuning strategies, interested readers can refer to the literatures [143, 144, 145, 146].
@@ III-C1 Fine Tuning
III-C2 Prompt Tuning
Prompt-based tuning initially introduced in NLP, offers an efficient method for adapting foundation models to downstream language tasks. This approach diverges from the traditional ”pre-training, fine-tuning” paradigm, which initializes and optimizes the weight parameters of pre-trained models using task-specific loss functions. Instead, prompt-based learning employs textual prompts to reformulate various downstream tasks within the framework of the original pre-trained task. Drawing inspiration from these NLP techniques, prompt tuning has subsequently been applied to the field of computer vision. In this context, vision prompt tuning can be categorized into three distinct approaches: vision-driven prompts, language-driven prompts, and vision-language prompts, as shown in Fig. 10.
@@ III-C5 Remapping Tuning
IV Key Advances of Recent Geoscience Foundation Models
With the applications of large foundation models for general purposes, the explorations of foundation models for remote sensing-based geoscience tasks have drawn a lot of attention from the research community. In this subsection, we will review the recent advances in GFMs, including the summary of existing foundation models from the view of large language models (LLMs), large vision models (LVMs), and large vision-language models (LVLMs). Corresponding LLMs, LVMs, and LVLMs for geoscience tasks are also summarized in Table II, Table III, and Table IV, respectively.
IV-A Large Language Models
LLMs have have emerged as a focal point of research in the domains of natural language processing (NLP) and computer vision. These models, characterized by their extensive transformer-based architectures, are designed to facilitate comprehensive understanding of both visual and textual data. LLMs have demonstrated unprecedented efficacy across a spectrum of language-related tasks, including but not limited to language modeling, text generation, and question answering [180, 97].
IV-A1 Basic LLMs
@@
Various BERT model variants have been developed to enhance its performance and efficiency. RoBERTa [107], developed by Facebook AI, improved BERT through more iterations and data augmentation techniques like dynamic masking and noising. ALBERT [108] introduced factorized embedding parameterization, cross-layer parameter sharing, and a sentence order prediction task to reduce model size and improve training speed. MacBERT [183] proposed replacing the [MASK] token with synonyms at mask locations instead of using the [MASK] tag. These variants demonstrate ongoing efforts to optimize BERT’s architecture, training process, and performance across various natural language processing tasks, showcasing the model’s adaptability and potential for further improvements in language understanding and generation.
@@
IV-A2 Advanced LLMs for Geoscience
@@
For Chinese geological texts, the authors in [190] developed a large-scale Chinese geological corpus (GeoCorpus) and a Chinese pretrained language model (CnGeoPLM) for geological domain, demonstrating that CnGeoPLM significantly outperforms the generic BERT model in geological NLP tasks like entity recognition, relationship extraction, and entity clustering. Wei et al. [191] introduced GeoBERTSegmenter, a GeoBERT-based Chinese word segmentation model tailored for geological texts, combining BERT, bidirectional recurrent neural network, and conditional random field to effectively address the segmentation of unregistered geological terms and common words, demonstrating superior performance in precision, recall, and F1 score compared to other methods.
@@ IV-A3 Applications and Disscussions
IV-B Large Vision Models
IV-B1 Basic Large Vision Models
@@
Figure 14: The Segment Anything Model (SAM) overview [213].
IV-B2 Advanced Large Vision Models for Geoscience
LVMs have been widely adopted in remote sensing and geoscience for various tasks, including remote sensing image analysis, satellite image classification, and climate and weather modeling. Most LVMs are based on the transformer architecture, which has proven effective in handling large-scale remote sensing data. Several studies have introduced general-purpose LVMs for remote sensing applications. For example, Sun et al. [18] introduced RingMo, a remote sensing foundation model framework using generative self-supervised learning, trained on two million remote sensing (RS) images to address domain gap and generalization issues. The work [19] proposed LVMs with a rotated varied-size window attention mechanism, demonstrating superior performance in detection, classification, and segmentation tasks. The authors in [214] explored the impact of increasing model parameters on performance, showing improved results and data efficiency in tasks like rotated object detection and semantic segmentation. Yao et al. [215] presented RingMo-Sense, a foundation model for spatiotemporal prediction in remote sensing, while Wang et al. [216] introduced RingMo-lite, a lightweight CNN-Transformer hybrid network optimized for edge devices.
@@ IV-B3 Applications and Discussions
LVMs have emerged as transformative technologies in geoscience and remote sensing, demonstrating remarkable capabilities across tasks such as image analysis, classification, segmentation, and climate modeling. The adoption of transformer-based architectures has enabled these models to effectively handle the complexity and scale of geospatial data, leading to significant improvements in performance and efficiency. Models like RingMo, Prithvi, ClimaX, and adaptations of the SAM have shown the potential to address domain-specific challenges in remote sensing, climate science, and Earth observation. As the field advances, key trends such as increased multimodal integration, enhanced spatial and temporal reasoning, improved explainability, and a focus on efficiency and ethical considerations are likely to shape future developments. These advancements promise to unlock new insights into our planet, enhance our ability to monitor and respond to environmental changes, and revolutionize how geoscientists, policymakers, and environmental managers analyze and understand complex geospatial data.
@@
IV-C Large Vision-Language Models
@@
IV-C1 Basic Large Vision-Language Models
@@
CLIP [142] is a prominent example that leverages a large-scale collection of image-text pairs from the Internet for contrastive learning, enabling the creation of a unified representation of multimodal data. Inspired by CLIP, various multimodal image FMs have been proposed, differing in their model structures [239, 240], feature representations [241, 242], multimodal feature fusion approaches [243], feature alignment loss functions [244], pretraining methods [245], and more. These advancements have significantly improved the performance of multimodal FMs.
@@
IV-C2 Advanced Large Vision-Language Models for Geoscience
@@
Evaluation and benchmarking of LVLMs in remote sensing have been crucial for understanding their capabilities and limitations. The authors in [254] conducted a comprehensive benchmark focusing on scene understanding, localization, counting, and change detection tasks, revealing that while these models excel in tasks like location understanding and captioning, their spatial reasoning limitations affect performance in object localization and counting. Li et al. [255] introduced VRSBench, a comprehensive vision-language benchmark for remote sensing image understanding, featuring a large dataset of verified captions, object references, and question-answer pairs.
@@
IV-C3 Applications and Disscussion
LVLMs have found diverse applications in remote sensing, including visual question answering [248], scene classification [261], object detection, change detection, image captioning, and multi-modal data fusion [237]. These applications have the potential to revolutionize fields such as urban planning, environmental monitoring, and disaster response. However, the advancement of LVLMs in remote sensing also presents challenges. While models like GeoChat [251] and lightweight architectures like LiT-4-RSVQA [249] improve accessibility and efficiency, issues of bias [250], interpretability, and ethical use remain critical concerns. The development of high-quality datasets [253] and benchmarks [255] is crucial for progress, as is the integration of domain-specific knowledge. As LVLMs continue to evolve, balancing their immense potential with careful consideration of fairness, privacy, and security will be essential to their successful application in Earth observation and geospatial analysis.
@@ IV-C3 Applications and Disscussion
V Challenges and Future Trends
@@ V-A Data Integration
V-A1 Variety of Data Sources
@@
V-A2 Data Quality and Compatibility
@@ V-B Model Complexity
V-B1 Multi-Scale Modeling
@@
V-B2 Simulating Complex Processes
@@ V-B4 Validation and Verification
V-B5 Model Interpretability
Interpretability remains an active field of research, continuously evolving with the development of new techniques. The degree of interpretability varies, as no model is entirely transparent or utterly opaque. Researchers are consistently working on methods to demystify the ”black box” nature of foundation models while recognizing their inherent limitations [29]. Foundation models, such as ChatGPT-4 and Claude-3.5, present interpretative challenges due to their substantial scale and complexity [16, 40]. Possessing billions of parameters, these models make it challenging to elucidate the reasons behind their specific outputs. Consequently, the complexity of GFMs can hinder stakeholders, including policymakers and the general public, from fully comprehending and utilizing the insights these models provide. This challenge may prevent the effective integration of scientific discoveries into decision-making processes crucial for environmental management and policy formulation.
V-C Uncertainty Quantification

### v5

@@ IV-A1 Vanilla Transformer
Figure 5: The Transformer model architecture [78].
IV-A2 Visual Transformer
@@
Figure 6: The framework of vision transformer (ViT) [90].
IV-A3 Vision-Language Transformer
@@ IV-B1 Supervised pre-training
IV-B2 Self-supervised pre-training
Self-supervised learning (SSL) has become a cornerstone in the development of foundation models (FMs), enabling them to acquire rich features and representations from vast amounts of unlabeled data [114], [115], [116]. This approach, often referred to as the ”dark matter” of intelligence, circumvents the limitations of supervised learning by leveraging a ”semiautomatic” process to derive ”labels” directly from the data, significantly reducing time and resource costs [115]. SSL has achieved remarkable success across various domains of deep learning, particularly in natural language processing with models such as BERT [97] and GPT-3 [117], and in computer vision with MAE [118] and DINOv2 [119]. The SSL workflow typically involves defining a pretext task based on unlabeled inputs to generate meaningful representations [114], [116], [120], [121]. These pretext tasks, which can be prediction-based, context-based, or generation-based, form the basis of the pre-training phase where the objective function is learned from unlabeled data [120]. The resulting representations are then transferred as initial weights for fine-tuning on specific downstream tasks. SSL methods can be broadly categorized into three approaches based on their pretext tasks: generative, contrastive, and predictive [114], [122]. This “pre-training, fine-tuning” paradigm has proven highly effective, allowing FMs to leverage the power of unlabeled data before being adapted to specific tasks with labeled data, often matching or surpassing purely supervised approaches in performance.
@@ IV-B2 Self-supervised pre-training
GANs [137] represent a powerful approach in generative methods, consisting of two competing networks: a generator $G:Z\rightarrow X$ that synthesizes fake samples from random noise in the latent space, and a discriminator $D:X\rightarrow[0,1]$ that attempts to distinguish between real and fake samples. This adversarial setup has proven effective for tasks such as image generation and super-resolution. Notable advancements include BigGAN [123], which introduces innovations sucha s a large-scale GAN architecture, conditional batch normalization, category conditioning, and orthogonal regularization, enabling the generation of high-quality, diverse, and category-specific images. In the domain of image enhancement, SRGAN [124] has demonstrated remarkable capabilities in super-resolution tasks, learning from a large dataset of downsampled images to recover high-resolution textures and effectively upscale low-resolution inputs. These applications showcase the potential of GANs not just for generating new images, but also for enhancing existing ones, pushing the boundaries of synthetic image generation and opening up numerous possibilities in image processing and computer vision tasks.
@@ IV-B3 Hybrid pre-training
IV-C Adaptation
Pre-trained GFMs can be adapted to specific tasks or datasets through fine-tuning, prompt tuning, adapter tuning, parameter tuning, and remapping tuning. For more tuning strategies, interested readers can refer to the literatures [143, 144, 145, 146].
@@ IV-C1 Fine-Tuning
IV-C2 Prompt Tuning
Prompt-based tuning initially introduced in NLP, offers an efficient method for adapting foundation models to downstream language tasks. This approach diverges from the traditional “pre-training, fine-tuning” paradigm, which initializes and optimizes the weight parameters of pre-trained models using task-specific loss functions. Instead, prompt-based learning employs textual prompts to reformulate various downstream tasks within the framework of the original pre-trained task. Drawing inspiration from these NLP techniques, prompt tuning has subsequently been applied to the field of computer vision. In this context, vision prompt tuning can be categorized into three distinct approaches, including vision-driven prompts, language-driven prompts, and vision-language prompts, as shown in Fig. 10.
@@ IV-C5 Remapping Tuning
V Key Advances of Recent Geoscience Foundation Models
With the applications of foundation models for general purposes, the explorations of foundation models for remote sensing-based geoscience tasks have drawn a lot of attention from the research community. In this subsection, we will review the recent advances in GFMs, including the summary of existing foundation models from the view of large language models (LLMs), large vision models (LVMs), large vision-language models (LVLMs), and foundation-model-based agents. Corresponding LLMs, LVMs, and LVLMs for geoscience tasks are also summarized in Table II, Table III, and Table IV, respectively.
V-A Large Language Models
LLMs have emerged as a focal point of research in the domains of natural language processing (NLP) and computer vision. These models, characterized by their extensive transformer-based architectures, are designed to facilitate comprehensive understanding of both visual and textual data. LLMs have demonstrated unprecedented efficacy across a spectrum of language-related tasks, including but not limited to language modeling, text generation, and question answering [180, 97].
V-A1 Basic LLMs
@@
Various BERT model variants have been developed to enhance its performance and efficiency. RoBERTa [107], developed by Facebook AI, improved BERT through more iterations and data augmentation techniques such as dynamic masking and noising. ALBERT [108] introduced factorized embedding parameterization, cross-layer parameter sharing, and a sentence order prediction task to reduce model size and improve training speed. MacBERT [183] proposed replacing the [MASK] token with synonyms at mask locations instead of using the [MASK] tag. These variants demonstrate ongoing efforts to optimize BERT’s architecture, training process, and performance across various natural language processing tasks, showcasing the model’s adaptability and potential for further improvements in language understanding and generation.
@@
V-A2 Advanced LLMs for Geoscience
@@
For Chinese geological texts, the authors in [190] developed a large-scale Chinese geological corpus (GeoCorpus) and a Chinese pretrained language model (CnGeoPLM) for geological domain, demonstrating that CnGeoPLM significantly outperforms the generic BERT model in geological NLP tasks such as entity recognition, relationship extraction, and entity clustering. Wei et al. [191] introduced GeoBERTSegmenter, a GeoBERT-based Chinese word segmentation model tailored for geological texts, combining BERT, bidirectional recurrent neural network, and conditional random field to effectively address the segmentation of unregistered geological terms and common words, demonstrating superior performance in precision, recall, and F1 score compared to other methods.
@@ V-A3 Applications and Disscussions
V-B Large Vision Models
V-B1 Basic Large Vision Models
@@
Figure 14: The Segment Anything Model (SAM) architecture [213].
V-B2 Advanced Large Vision Models for Geoscience
LVMs have been widely adopted in remote sensing and geoscience for various tasks, including remote sensing image analysis, satellite image classification, and climate and weather modeling. Most LVMs are based on the transformer architecture, which has proven effective in handling large-scale remote sensing data. Several studies have introduced general-purpose LVMs for remote sensing applications. For example, Sun et al. [18] introduced RingMo, a remote sensing foundation model framework using generative self-supervised learning, trained on two million remote sensing (RS) images to address domain gap and generalization issues. The work [19] proposed LVMs with a rotated varied-size window attention mechanism, demonstrating superior performance in detection, classification, and segmentation tasks. The authors in [214] explored the impact of increasing model parameters on performance, showing improved results and data efficiency in tasks such as rotated object detection and semantic segmentation. Yao et al. [215] presented RingMo-Sense, a foundation model for spatiotemporal prediction in remote sensing, while Wang et al. [216] introduced RingMo-lite, a lightweight CNN-Transformer hybrid network optimized for edge devices.
@@ V-B3 Applications and Discussions
LVMs have emerged as transformative technologies in geoscience and remote sensing, demonstrating remarkable capabilities across tasks such as image analysis, classification, segmentation, and climate modeling. The adoption of transformer-based architectures has enabled these models to effectively handle the complexity and scale of geospatial data, leading to significant improvements in performance and efficiency. Models such as RingMo, Prithvi, ClimaX, and adaptations of the SAM have shown the potential to address domain-specific challenges in remote sensing, climate science, and Earth observation. As the field advances, key trends such as increased multimodal integration, enhanced spatial and temporal reasoning, improved explainability, and a focus on efficiency and ethical considerations are likely to shape future developments. These advancements promise to unlock new insights into our planet, enhance our ability to monitor and respond to environmental changes, and revolutionize how geoscientists, policymakers, and environmental managers analyze and understand complex geospatial data.
@@
V-C Large Vision-Language Models
@@
V-C1 Basic Large Vision-Language Models
@@
CLIP [142] is a prominent example that leverages a large-scale collection of image-text pairs from the Internet for contrastive learning, enabling the creation of a unified representation of multimodal data. Inspired by CLIP, various multimodal image FMs have been proposed, differing in their model structures [239, 240], feature representations [241, 242], multimodal feature fusion approaches [243], feature alignment loss functions [244], pre-training methods [245], and more. These advancements have significantly improved the performance of multimodal FMs.
@@
V-C2 Advanced Large Vision-Language Models for Geoscience
@@
Evaluation and benchmarking of LVLMs in remote sensing have been crucial for understanding their capabilities and limitations. The authors in [254] conducted a comprehensive benchmark focusing on scene understanding, localization, counting, and change detection tasks, revealing that while these models excel in tasks such as location understanding and captioning, their spatial reasoning limitations affect performance in object localization and counting. Li et al. [255] introduced VRSBench, a comprehensive vision-language benchmark for remote sensing image understanding, featuring a large dataset of verified captions, object references, and question-answer pairs.
@@
V-C3 Applications and Disscussion
LVLMs have found diverse applications in remote sensing, including visual question answering [248], scene classification [261], object detection, change detection, image captioning, and multi-modal data fusion [237]. These applications have the potential to revolutionize fields such as urban planning, environmental monitoring, and disaster response. However, the advancement of LVLMs in remote sensing also presents challenges. While models such as GeoChat [251] and lightweight architectures such as LiT-4-RSVQA [249] improve accessibility and efficiency, issues of bias [250], interpretability, and ethical use remain critical concerns. The development of high-quality datasets [253] and benchmarks [255] is crucial for progress, as is the integration of domain-specific knowledge. As LVLMs continue to evolve, balancing their immense potential with careful consideration of fairness, privacy, and security will be essential to their successful application in Earth observation and geospatial analysis.
@@ V-C3 Applications and Disscussion
V-D Foundation-Model-based Agents
The concept of ”agents” has a rich and multifaceted history, traversing diverse fields from philosophy [269] and cognitive science to artificial intelligence [270]. Originally rooted in discussions of agency and autonomy in human behavior, the term has evolved to encompass a wide range of entities capable of perceiving their environment and taking actions to achieve specific goals. For more detials about general foundation-model-based agents, please refer to [271, 272].
V-D1 Basic Foundation-Model-based Agents
As shown in Fig. 16, a typical large language model (LLM)-based agent consists of three main components: brain, perception, and action.
  • Brain: The brain module is responsible for basic tasks, including memorizing, thinking, and decision-making.
  • Perception: The perception module processes the input data, such as text, images, or other modalities, and converts it into a format that the brain can understand.
  • Action: The the action module carries out the execution using tools and influences the surroundings.
When asked about rain, the agent’s perception module interprets the question, while the brain module reasons based on current weather conditions and online weather reports. The action module then responds and potentially offers an umbrella, with this process repeating to allow continuous interaction and feedback between the agent and its environment.
Figure 16: Conceptual framework of LLM-based agent [271].
V-D2 Advanced Foundation-Model-based Agents for Geoscience
There are only several studies on foundation-model-based agents for geoscience, particularly remote sensing (RS). For example, RS-ChatGPT [273] includes the process of prompt template generation, task planning, task execution, and response generation, as shown in Fig. 17. RS-ChatGPT operates through a three-step workflow: prompt template generation, task planning, and task execution with response generation. Initially, it creates a prompt template that guides ChatGPT in processing remote sensing tasks, incorporating strict file handling, tool-based execution, and image captioning. Next, it plans tasks by generating detailed descriptions of various remote sensing functions, which are integrated into the prompt template to enhance ChatGPT’s understanding. Finally, ChatGPT executes the planned tasks by selecting and applying appropriate tools to preprocess remote sensing images, iteratively feeding results back as new observations until a final response is generated based on all executed tasks. This workflow enables RS-ChatGPT to effectively analyze remote sensing data despite ChatGPT’s inability to directly access images, with the potential for further expansion and improvement of its capabilities.
Figure 17: The workflow of Remote Sensing ChatGPT [273].
Similarly, the authors in [274] proposed an interactive Change-Agent system, innovating RS image change interpretation by combining a multilevel change interpretation (MCI) model with a large language model. This integration allows for comprehensive analysis of Earth’s surface changes, including detection, captioning, object counting, and cause analysis. The MCI model, trained on a new dataset, uses pixel-level and semantic-level processing enhanced by a novel interaction layer. Experiments show top performance in both change detection and description. Xu et al. [275] explored RS-Agent, which is an advanced RS AI that combines a large language model, specialized image processing tools, and expert knowledge. This integration allows it to understand complex queries, perform multi-step analyzes, and provide professional insights in RS tasks. Tested on various datasets, RS-Agent shows superior performance in scene classification, visual question answering, and object counting, demonstrating its ability to handle complex RS applications more effectively than previous models.
V-D3 Applications and Discussions
The development of foundation-model-based agents for geoscience, particularly RS, has the potential to revolutionize the field by enabling intelligent systems capable of understanding and processing complex geospatial data. RS-ChatGPT [273], Change-Agent [274], and RS-Agent [275] represent a significant step forward in this direction, demonstrating the feasibility of using LLMs to perform a wide range of RS tasks, including scene classification, land use classification, object detection, image captioning, edge detection, polygonization, object counting, and change detection. In the future, foundation-model-based agents are expected to play an increasingly important role in geoscience research, enabling more efficient data analysis, decision-making, and knowledge discovery.
VI Challenges and Future Trends
@@ VI-A Data Integration
VI-A1 Variety of Data Sources
@@
VI-A2 Data Quality and Compatibility
@@ VI-B Model Complexity
VI-B1 Multi-Scale Modeling
@@
VI-B2 Simulating Complex Processes
@@ VI-B4 Validation and Verification
VI-B5 Model Interpretability
Interpretability remains an active field of research, continuously evolving with the development of new techniques. The degree of interpretability varies, as no model is entirely transparent or utterly opaque. Researchers are consistently working on methods to demystify the ”black box” nature of foundation models while recognizing their inherent limitations [29]. Foundation models, such as ChatGPT-4 and Claude-3.5, present interpretative challenges due to their substantial scale and complexity [16, 40]. Possessing billions of parameters, these models make it challenging to elucidate the reasons behind their specific outputs. Consequently, the complexity of GFMs can hinder stakeholders, including policymakers and the public, from fully comprehending and utilizing the insights these models provide. This challenge may prevent the effective integration of scientific discoveries into decision-making processes crucial for environmental management and policy formulation.
VI-C Uncertainty Quantification